In 1968 the Micron Company stock went up from $26 to $32 per share and

      paid a $1.50 dividend. Inflation that year increase from 12% to 15% for the

      year. What was the real total return for Micron shareholders in 1968?
Answer #1

Nominal Return of Dividend = (32-26+1.50)/26 = 28.8462%
Inflation = 15%

Total Real Return in 1968 = (1+Nominal Rate)/(1+Inflation) -1 = (1+28.8462%)/(1+15%)-1 = 12.04%

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well